OpenStudy (anonymous):

The probability that James will win two races in the next two days is 1 over 8 for tomorrow and 1 over 10 for the day after tomorrow. What is P(win tomorrow, then win the day after tomorrow)?

OpenStudy (anonymous):

multiply

OpenStudy (anonymous):

at least that is what i think they want you to do looks like it was worded by a 12 year old also it assumes the events are independent which is rather unlikely in any case i am almost sure they want you to compute \[\frac{1}{8}\times \frac{1}{10}\]

OpenStudy (anonymous):

so 80